The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
16/100

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

16 / 100 = (4 × 4) / (25 × 4)
= 4 / 25

Thanks